Kenosha Winter Services is a small, family-owned snow-removal business. For its services, the company has always charged a flThe total cost of operating the company for the year is $532,500, which includes the following costs: Wages Supplies Snow remResource consumption is distributed across the activities as follows: Distribution of Resource Consumption across Activity CoRequired 1. Prepare the first-stage allocation of costs to the activity cost pools. (Do not leave any empty spaces; input a 02. Compute the activity rates for the activity cost pools. (Round your answers to 2 decimal places.) Activity Cost Pool Snow3. The company recently completed a 5,400-square-metre snow removal job at Hometown Hardware-a 94-kilometre round-trip journe4. The revenue from the Hometown Hardware job was $685.80 (5,400 square metres at $12.70 per hundred square metres). Prepare
Answer #1
1
Snow removal Travel to Jobs Job support Other Total
Wages 144900 41400 0 20700 207000
Supplies 49500 0 0 0 49500
Snow removal equipment depr. 23600 0 0 5900 29500
Vehicle expenses 0 34650 0 14850 49500
Office expense 0 0 31600 47400 79000
President’s compensation 0 0 53100 64900 118000
Total cost 218000 76050 84700 153750 532500
2
Activity Cost Pool Activity Rate   
Snow removal 5.00 per hundred square metres =218000/43600
Travel to jobs 3.00 per kilometre =76050/25350
Job support 423.50 per job =84700/200
3
Activity Cost Pool ABC cost
Snow removal 270.00 =54*5
Travel to jobs 282.00 =94*3
Job support 423.50
Total 975.50
4
Sales 685.80
Costs:
Snow removal 270.00
Travel to jobs 282.00
Job support 423.50 975.50
Product margin -289.70